Use cylindrical shells to find the volume of the solid that results when the region enclosed by x = y^2 and x = y is revolved about the line y= -1
1
Expert's answer
2020-09-07T17:25:33-0400

Horizontal axis of revolution


"V=2\pi\displaystyle\int_{c}^dp(y)h(y)dy"

"y^2=y=>y_1=0, y_2=1"

"p(y)=y-(-1)=y+1"

"h(y)=y-y^2"


"V=2\pi\displaystyle\int_{0}^1(y+1)(y-y^2)dy="

"=2\pi\displaystyle\int_{0}^1(y^2-y^3+y-y^2)dy="

"=2\pi\big[\dfrac{y^2}{2}-\dfrac{y^4}{4}\big]\begin{matrix}
 1 \\
 0
\end{matrix}=\dfrac{1}{2}\pi\ (units^3)"